Susan decides to deposit $50,000 into a bank that offers 4.75% interest compounded monthly. After six years, what will Susan's account total?
Answer #1
Annual Interest = 4..75%
Monthly interest = 4.75/12 = 0.3958%
Effective rate = (1.003958)^12 - 1 =4.85%
Amount deposited 50000
Multiply: FVF at 4.85% for 6th yr 1.32865
Amount in the account after 6yrs 66432.5
Answer s 66432.50